OpenStudy (anonymous):

Solve the problem. Show all work. The area of a triangle with basse b and heeight h is given by the formula A=1/2bh Find the area of a triangle with base 2 meters and height 12 meters.

OpenStudy (anonymous):

1/2(2*12)=1/2(24)=12